# ReAgent: Reversible Multi-Agent Reasoning for Knowledge-Enhanced Multi-Hop QA

### v1

@@ 3.5 Conflict Management and Backtracking
This part formalizes how the system detects contradictions and reverts to consistent states. At each time $t$, the set of all accepted assertions is $\Phi^{t}\;=\;\bigcup_{i=1}^{n}\Phi_{i}^{t}.$ A conflict occurs if $\Phi^{t}$ entails both $\phi$ and $\neg\phi$ for some proposition $\phi$, meaning $\mathrm{SAT}(\Phi^{t})$ fails. To resolve the issue, the system follows a two-level approach:
Local Backtracking.
Each agent $A_{i}$ maintains a backtracking graph $\mathrm{LBG}_{i}$, which logs states $\Phi_{i}^{r}$ at selected checkpoints $r$. If $\mathrm{Conflict}_{i}(\Phi_{i}^{t})$ indicates an internal contradiction, $A_{i}$ tries to revert to a prior consistent node $r<t$:
$\mathrm{BacktrackLocal}(A_{i},\,r):\quad\Phi_{i}^{t}\;\longrightarrow\;\Phi_{i}^{r}.$
After local rollback, the agent re-evaluates newly arriving evidence.
Global Conflict and Holistic Updates.
When contradictions span multiple agents, the Supervisor Agent identifies a minimum conflict set of assertions that must be revised. A global backtracking operation
$\mathrm{BacktrackGlobal}\bigl{(}\sigma^{t},\,r\bigr{)}$
reverts all agents from time $t$ to $r$. The system discards any statements introduced between $r+1$ and $t$, restoring consistency. If the conflict cannot be removed even after a global rollback, the Controller might enforce strategic overrides or disclaim an answer.
@@
Zero-shot
Qwen-2.5-14b-Instruct 0.205 0.316 0.258 0.308 0.042 0.131 0.168 0.251
Qwen-2.5-32b-Instruct 0.191 0.304 0.190 0.294 0.041 0.133 0.140 0.243
Qwen-2.5-72b-Instruct 0.238 0.355 0.286 0.338 0.049 0.141 0.191 0.278
Llama-3.3-70b-Instruct 0.280 0.415 0.290 0.349 0.063 0.155 0.211 0.306
DeepSeek-v3 0.293 0.410 0.258 0.332 0.042 0.145 0.197 0.295
Gemini-1.5-flash 0.236 0.348 0.267 0.311 0.055 0.148 0.186 0.269
Gemini-2.0-flash 0.317 0.434 0.335 0.391 0.075 0.170 0.242 0.331
GPT-4o 0.359 0.497 0.346 0.417 0.100 0.217 0.268 0.377
Knowledge Augmented
Qwen-2.5-14b-Instruct 0.594 0.768 0.623 0.725 0.305 0.453 0.507 0.648
Qwen-2.5-32b-Instruct 0.576 0.756 0.537 0.694 0.324 0.460 0.479 0.636
Qwen-2.5-70b-Instruct 0.604 0.772 0.664 0.753 0.326 0.475 0.531 0.667
Llama-3.3-70b-Instruct 0.624 0.790 0.698 0.783 0.332 0.470 0.551 0.681
DeepSeek-v3 0.620 0.792 0.673 0.769 0.368 0.505 0.553 0.688
Gemini-1.5-flash 0.622 0.784 0.644 0.732 0.314 0.435 0.526 0.680
Gemini-2.0-flash 0.623 0.790 0.638 0.744 0.306 0.441 0.522 0.685
GPT-4o 0.612 0.794 0.677 0.780 0.356 0.506 0.548 0.693
@@
Table 1: Performance of different models across three multi-hop QA datasets. In each column, the highest performance is highlighted in bold; and within each method group, the top performer is underlined.
4 Experiments
4.1 Experimental Setup
We assessed ReAgent through three widely used knowledge-intensive multi-hop QA datasets: HotpotQA (Yang et al., 2018), 2WikiMultiHotQA (Ho et al., 2020) and Musique (Trivedi et al., 2022). We show how ReAgent produces better results compared to baseline models and effectively solves problems requiring continuously corrected reasoning.
Datasets
HotpotQA is a large-scaled, open-domain dataset encouraging multi-hop reasoning. The dataset provides full Wikipedia passages for information retrieval and integration. 2WikiMultiHopQA selects distinct Wikipedia domains for each question, focusing on evaluating the model’s multi-hop inference over different resources rather than one document. Musique is a challenging dataset requiring model’s ability to reason over multiple dispread sentences.
Following previous work Trivedi et al. (2023); Press et al. (2023); Gutiérrez et al. (2024), we utilize 1000 random samples from each validation set and corresponding texts as the knowledge base.
Baselines.
We include both proprietary and open-sourced LLMs with multiple sizes for evaluation: Qwen-2.5 Yang et al. (2024a) of varying parameter scales, the newly Llama-3.3 Meta (2024) and DeepSeek-v3 Liu et al. (2024), Gemini Team et al. (2023) variants (Gemini-1.5-flash-2024-09, Gemini-2.0-flash-2025-02 as well as GPT-4o-2024-08-06 Hurst et al. (2024).
Settings.
We hold two distinct settings for the performance evaluation: Zero-shot with CoT only provides the question, expecting to validate LLM’s performance with internal knowledge; Knowledge-Augmented with CoT Liang et al. (2024); Gutiérrez et al. (2024) setting introduces related documents for reasoning, evaluating model’s capability of information integration especially processing multiple mixed evidence.
Implementation.
For super large-size LLMs (DeepSeek-v3, Gemini families, GPT-4o), we access through API, for medium-size LLMs, we access from their official repositories. We set the temperature as $0.2$ to get deterministic answers as well as somehow maintain flexibility. We utilize the default settings of other parameters. ReAgent utilizes GPT-4o as the backbone for execution. Specifically, we set the temperature for decomposition agent as $0.8$, the remaining as $0.6$. All prompts are available in Appendix A.
Metrics.
We measure Exact Match (EM) and F1 scores for the QA evaluation.
@@ 4.2 Results and Analysis
Table 1 illustrates the comprehensive performance. ReAgent outperforms all the baseline models in both metrics across all datasets. Specifically, our framework achieves an average EM of 0.571 and an average F1 of 0.701, outperforming GPT-4o by 2.3% in EM and 0.8% in F1.
External Knowledge Matters.
Definitely, we observe that zero-shot performance is substantially lower, reflecting the difficulty of multi-hop reasoning without relevant background. With intensive knowledge, all models outperform their zero-shot versions in different levels, presenting their internal reasoning abilities.
Scaling-law not always works.
We carefully select baseline models across small(14b), medium(70b), large(>100b) size for comparisons. Interestingly, the results show that smaller models sometimes outperform larger models. For example, Qwen-2.5-14b model generally defeats its 32b counterparts in both zero-shot and knowledge augmented settings. Similarly, Llama-3.3 approach or even surpass the super large-scale models (e.g. Gemini-1.5-flash and DeepSeek-v3 in zero-shot, Gemini-2.0-flash and GPT-4o in knowledge-intensive setting). This suggests that multi-hop QA can be aided by carefully tuned architectures, especially if they incorporate chain-of-thought or knowledge-aware strategies.
Overall, we emphasized the value of reversible reasoning mechanisms to mitigate error propagation. The data in Table 1 align with this assumption: even with strong models, single-pass reasoning can fail unless the correct context is identified or re-checked. Our method’s ability to backtrack helps resolve contradictions, leading to more stable performance in multi-hop scenarios.
Figure 4: Ablation Study Results Comparison.
@@ 4.3 Ablation Study
To gauge the importance of the backtracking mechanism, we conduct an ablation study by disabling it under the same settings. Specifically, the system then proceeds in a strictly forward fashion, without the ability to revert to earlier intermediate states. Fig.4 compares performance between different backbones, with and without backtracking, on three datasets. As shown, the performance drop highlights the importance of backtracking in mitigating error propagation, where an early misstep without backtracking cannot be corrected, resulting in deteriorated performance. Notably, the DeepSeek-v3 backbone exhibits general improvements with backtracking, demonstrating its robustness across various settings.
@@ 5 Case Study
We illustrate our multi-agent backtracking (ReAgent) through a single-culprit puzzle with four suspects {A, B, C, D}. Each suspect gives statements about who might be guilty or lying. The puzzle’s only correct solution is that B is the culprit, but identifying this requires partially retracting initial assumptions along the way. Detailed puzzle statement can be found in Appendix B.
Why a Rollback Mechanism is Needed.
Naive single-pass (or single-thread) models often fixate prematurely on one suspect, disregard contradictory evidence, and produce unsound conclusions. By contrast, our multi-agent approach enumerates possible culprits in parallel, detects conflicts, and rolls back to revise incorrect assumptions.
Below is a simplified agentic workflow capturing the time steps $T_{1}$ through $T_{7}$ during which each agent processes statements, detects a conflict, and reverts.
Algorithm 2 ReAgent Multi-Agent Puzzle Solving
1:Input: Puzzle statements from A, B, C, D.
3:Initialization
[…]
6:Round 1: Checking A as culprit
10:Round 2: Checking B as culprit
13:$T_{6}:$ Conflict Detector sees consistency, no rollback needed.
14:Rounds 3 and 4: Checking C or D as culprit
Discussion.
This staged procedure allows each culprit hypothesis to be explored (lines 7–13 of Algorithm 2). Any contradictions are escalated at $T_{2}$ or $T_{7}$ to the Conflict Detector, which triggers a rollback to the prior consistent node. Ultimately, only the assignment “$B$ is culprit” withstands these iterative checks. Without rollback, a linear solver often locks in a wrong suspect early and fails to correct itself. For extended discussions of single-thread LLM failures, and the complete multi-agent transcripts, see Appendix B and  C.
6 Related Work
6.1 Prompting and Iterative Reasoning
Large Language Models (LLMs) can solve complex tasks by leveraging prompting strategies that encourage step-by-step reasoning, often referred to as chain-of-thought (CoT) prompting Wei et al. (2022), which has demonstrated improved performance in arithmetic, commonsense, and symbolic reasoning tasks, particularly in sufficiently large models. Sampling multiple reasoning paths and selecting the most consistent answer, known as self-consistency, further boosts accuracy by treating final answers as a “majority vote” across different solutions Wang et al. (2023). To reduce the need for hand-tuned prompts, automated prompt engineering techniques have been explored. One approach, Automatic Prompt Engineer (APE), treats prompt generation as a search problem: one LLM proposes candidate instructions while another evaluates them, allowing the system to refine prompts over multiple tasks Zhou et al. (2023). LLMs can thus act as both reasoning engines and prompt designers. Beyond one-shot responses, iterative refinement frameworks empower LLMs to correct and enhance their outputs. For instance, the ReAct framework interleaves reasoning traces with actions (e.g., external API calls), enabling LLMs to incorporate feedback and reduce factual errors Yao et al. (2023). Other methods like Reflexion Shinn et al. (2023) and Self-Refine Madaan et al. (2023) introduce self-feedback loops, where a model critiques its own response and revises it in subsequent attempts, leading to more reliable and transparent solutions. Prompting methods have also expanded into multimodal settings, where models process both text and visual inputs. In Multimodal-CoT, the system first generates a textual chain-of-thought integrating visual information, then produces an answer from that combined rationale Zhang et al. (2023).
6.2 Multi-Agent Collaboration, Debate, and Scalability
Recent work explores multi-agent approaches, where multiple LLMs (or instances of the same model) collaborate or compete toward a shared goal. For instance, the CAMEL framework assigns different roles (e.g., “user” and “assistant”) to LLM agents that communicate in a dialogue-based fashion, breaking down tasks more systematically Li et al. (2023). Role specialization allows agents to focus on different aspects of the problem, and voting or consensus mechanisms—akin to self-consistency—can yield more robust answers. Debate-oriented methods push multi-agent interactions further by having agents argue for different viewpoints. After hearing opposing arguments, a human or model judge selects the most convincing solution Khan et al. (2024). Similarly, in an interconsistency framework, multiple agents provide arguments and counterarguments over multiple rounds, eventually reaching a consensus Xiong et al. (2023). Such adversarial or critical exchanges can improve truthfulness and correctness, though an overly dominant model can bias outcomes unless moderated by a sufficiently capable judge Xiong et al. (2023). These multi-agent and multimodal systems introduce scalability and coordination challenges. As more agents or steps are added, communication overhead can grow significantly, making real-time collaboration difficult. Research highlights the need for efficient structures—e.g., hierarchical controllers or learned protocols—to balance the benefits of specialized agents with the system’s complexity Li et al. (2023); Xiong et al. (2023).

### v2

@@ 4.4 Conflict Management and Backtracking
This part formalizes how the system detects contradictions and reverts to consistent states. At each time $t$, the set of all accepted assertions is $\Phi^{t}$. A conflict occurs if $\Phi^{t}$ entails both $\phi$ and $\neg\phi$ for some proposition $\phi$, meaning $\mathrm{SAT}(\Phi^{t})$ fails. To resolve the issue, the system follows a two-level approach:
Local Backtracking. Each agent $A_{i}$ maintains a backtracking graph $\mathrm{LBG}_{i}$, which logs states $\Phi_{i}^{r}$ at selected checkpoints $r$. If an internal contradiction $\mathrm{Conflict}_{i}(\Phi_{i}^{t})$ is detected by $A_{V}$, local backtracking is performed to revert $A_{i}$ to a prior consistent node $r<t$:
$\mathrm{BacktrackLocal}(A_{i},\,r):\Phi_{i}^{t}\longrightarrow\Phi_{i}^{r}.$
After the local rollback, the agent re-evaluates newly arriving evidence.
Global Backtracking. When contradictions span multiple agents, the $A_{S}$ identifies a minimum conflict set of assertions that must be revised. A global backtracking operation: $\mathrm{BacktrackGlobal}\bigl{(}\sigma^{t},\,r\bigr{)}$ reverts all agents from time $t$ to $r$. The system discards any statements introduced between $r+1$ and $t$, restoring consistency. If the conflict cannot be removed even after a global rollback, the Controller might enforce strategic overrides or disclaim an answer. The corresponding algorithm is detailed in Appendix E.
5 Experiments
5.1 Experimental Setup
We evaluate ReAgent on three widely used, knowledge-intensive multi-hop QA benchmarks: HotpotQA (Yang et al., 2018), 2WikiMultiHopQA (Ho et al., 2020), and MuSiQue (Trivedi et al., 2022). We compare ReAgent against three main groups of baselines: (1) standard LLMs, (2) dedicated reasoning models, and (3) agent-based models. Experimental results demonstrate that ReAgent consistently outperforms all baseline groups and is particularly effective in solving tasks that require iteratively corrected reasoning.
Datasets. HotpotQA is a large-scale, open-domain dataset that explicitly promotes multi-hop reasoning by requiring the integration of information across multiple full-length Wikipedia passages. 2WikiMultiHopQA selects distinct Wikipedia domains for each question, focusing on evaluating the model’s multi-hop inference over different resources rather than one document. Musique is a challenging dataset requiring model’s ability to reason over multiple dispread sentences. Following previous work (Trivedi et al., 2023; Press et al., 2023; Gutiérrez et al., 2024), we utilize 1000 random samples from each validation set and corresponding texts as the knowledge base.
Baselines. We meticulously categorize diverse models into different groups to compare their capabilities in multi-hop QA. (1) Regular Models: We select non-reasoning models in this group. The models include: Llama-4  [^2] https://ai.meta.com/blog/llama-4-multimodal-intelligence/, Qwen-2.5 (Yang et al., 2024a) series, DeepSeek-V3-2024-03 (Liu et al., 2024), Genmini-1.5-Flash-2024-09 (Gemini et al., 2023), Gemini-2.0-flash-2025-02 (Gemini et al., 2023), GPT-4o-latest (Hurst et al., 2024), GPT-4.1 (Meta AI, 2024), GPT-4o with CoT (Wei et al., 2023). (2) Reasoning Models: This group includes several strong reasoning baselines for comparison. The models include: Qwen-3-Thinking (Qwen3, 2025) in 32B and 253B size, DeepSeek-R1 (DeepSeek-AI, 2025), Gemini-2.5-pro (Google DeepMind, 2025), O1 (OpenAI, 2024), and O3 (OpenAI, 2025). (3) Agentic Models. Chain-of-Agents (CoA) (Zhang et al., 2024), HippRAG (Gutiérrez et al., 2025), KAG (Liang et al., 2024), and our method. To compare fairly, we employ GPT-4o as the backbone for both COT and Agentic Models.
Implementation. For large-scale LLMs (e.g., DeepSeek-V3, the Gemini family, GPT-4o, and Qwen-3-235B), we access the models via API. For the medium-sized LLMs, we use their official open-source repositories. The temperature is set to $0.3$ to ensure deterministic outputs, while other parameters follow their default settings. ReAgent is implemented using GPT-4o as the backbone model. Specifically, we set the temperature to $0.8$ for the decomposition agent to encourage diverse reasoning paths, and $0.6$ for all other agents. The cost analysis for proprietary models and agentic methods is presented in Appendix B. For open-source models, all experiments are conducted using four A100-80G GPUs.
Metrics. We measure Exact Match (EM) and F1 scores for the QA evaluation.
@@ 5.1 Experimental Setup
Regular Models
Llama-4-Scout-17B-16E-Instruct 0.263 0.389 0.332 0.467 0.107 0.185 0.234 0.346
DeepSeek-V3 0.352 0.491 0.466 0.579 0.223 0.33 0.347 0.467
Qwen-2.5-32B-Instruct 0.372 0.509 0.557 0.663 0.159 0.273 0.363 0.482
Qwen-2.5-72B-Instruct 0.363 0.519 0.543 0.631 0.222 0.327 0.376 0.492
Gemini-1.5-Flash 0.374 0.488 0.563 0.650 0.208 0.310 0.381 0.482
Gemini-2.0-Flash 0.371 0.490 0.538 0.651 0.246 0.338 0.385 0.493
GPT-4o 0.381 0.549 0.517 0.649 0.245 0.379 0.381 0.525
GPT-4.1 0.389 0.563 0.544 0.665 0.271 0.413 0.401 0.547
CoT (GPT-4o) 0.408 0.531 0.558 0.638 0.272 0.360 0.413 0.509
Reasoning Models
Qwen-3-32B-Thinking 0.332 0.474 0.241 0.357 0.387 0.511 0.387 0.511
DeepSeek-R1 0.356 0.483 0.601 0.707 0.298 0.416 0.418 0.535
Qwen-3-235B-A22B-Thinking 0.361 0.506 0.624 0.729 0.271 0.387 0.418 0.540
Gemini-2.5-Pro 0.430 0.560 0.743 0.829 0.383 0.491 0.518 0.626
O1 0.505 0.661 0.656 0.758 0.417 0.551 0.526 0.656
O3 0.535 0.696 0.706 0.787 0.442 0.579 0.561 0.687
Agentic Models (w.GPT-4o)
CoA Zhang et al. (2024) 0.391 0.558 0.575 0.697 0.239 0.361 0.402 0.539
HippoRAG (Gutiérrez et al., 2025) 0.528 0.717 0.633 0.725 0.353 0.507 0.504 0.649
KAG (Liang et al., 2024) 0.603 0.782 0.681 0.781 0.348 0.489 0.544 0.684
@@
Table 1: Performance of different models across three multi-hop QA datasets. In each column, the highest and the second highest performance is highlighted in red and blue; and within each method group, the top performer is underlined.
@@ 5.2 Results and Analysis
Table 1 presents the overall performance across all datasets. ReAgent consistently outperforms all baseline models on both EM and F1 metrics. Specifically, it achieves an average EM of 0.571 and an average F1 of 0.701, surpassing the strongest baseline—knowledge-augmented GPT-4o—by 2.3% in EM and 0.8% in F1. It also outperforms strong reasoning models, including O1 and O3. Moreover, it also surpasses the recent agentic models such as CoA, HippoRAG, and KAG across all datasets. We summarize the following insights:
Agent-based and reasoning enhances performance. Some reasoning models and agentic models have a competitive performance (such as O1, O3, and KAG) than regular models. Notably, ReAgent outperforms O3, one of the strongest reasoning models, and improves significantly over GPT-4o+CoT in both EM and F1. This suggests that modular execution and collaboration among agents provide a clearer advantage in complex, multi-hop settings. While ReAgent can theoretically use stronger backbones like O1 or O3, their high cost (645\$ and 546\$, respectively, as shown in Appendix B) makes them not a good option. Our focus is to show that even with GPT-4o, ReAgent is able to beat the expensive O1 and O3.
ReAgent excels in manageable context length. Specifically, ReAgent achieves the highest F1 score on HotpotQA (0.795), demonstrating strong performance in tasks that require structured reasoning over relatively short contexts. In contrast, Gemini-2.5-Pro performs well on 2Wiki but underperforms on other datasets, while O1 and O3 show stronger results on Musique but fall short on HotpotQA. This discrepancy is partly attributed to the increased complexity of the Musique dataset, which involves longer contexts. While ReAgent is built on GPT-4o and may be less optimized for long documents than dedicated reasoning models, its agentic design, including traceback and self-check mechanisms, proves particularly effective in scenarios requiring precision, stepwise planning, and robust verification. These strengths enable ReAgent to outperform other models on tasks like HotpotQA, highlighting its superior general reasoning capabilities within manageable context lengths.
Overall, we emphasized the value of reversible reasoning mechanisms to mitigate error propagation. Results in Table 1 align with this assumption: even with strong models, single-pass reasoning can fail unless the correct context is identified or re-checked. Our method’s ability to backtrack helps resolve contradictions, leading to more stable performance in multi-hop scenarios.
@@ 5.3 Ablation Study
Effectiveness Analysis of Backtracking Mechanism. We conduct an ablation study by disabling backtracking under the same settings to verify its importance. In this setup, the system operates strictly forward without the ability to revert to earlier intermediate states. Figure 3 compares performance between different backbones, with and without backtracking, on HotpotQA dataset. The performance drop highlights the importance of backtracking in mitigating error propagation, where an early misstep without backtracking cannot be corrected, resulting in deteriorated performance. Notably, the DeepSeek-V3 backbone exhibits general improvements with backtracking, demonstrating its robustness across various settings.
Impact of Different Local Backtracking (BT) Depth. To further analyze the impact of local backtracking depths (the number of steps the system is allowed to revert), we analyze the performance of our proposed ReAgent under different settings, as shown in Figure 4 (left). The results show that on two selected backbones, DeepSeek-V3 and GPT-4o, the performance of ReAgent improves as the local backtracking depth increases, suggesting that a deeper backtracking depth benefits the agent by enabling it to effectively recover from earlier reasoning errors. However, the benefits gradually saturate, indicating that further increasing the backtracking depth yields diminishing returns.
Figure 3: Ablation Study on Local and Global Backtracking (BT): EM comparison on HotpotQA using GPT-4o and DeepSeek-V3.
Impact of the Number of Decomposed Sub-Questions. Figure 4 (right) demonstrates how a different number of decomposed sub-questions affects performance. The results show that decomposing the input question into multiple sub-questions significantly improves performance. Specifically, using 3 sub-questions is the optimal trade-off choice, with an improvement of 0.161 EM on the GPT-4o and 0.142 EM on DeepSeek-V3 compared to directly processing a single question. Beyond this point, increasing the number of sub-questions brings only marginal gains but higher costs.
Figure 4: Ablation Study on Backtracking Depth (left) and Number of Decomposed Sub-Questions (right): EM comparison on HotpotQA using GPT-4o and DeepSeek-V3.
Figure 5: Case study comparing GPT-O3 (left), CoA (middle), and ReAgent (right) on HotpotQA. The back-tracking mechanism enhances iterative reasoning, enabling conflict detection and correction to reach the correct answer.
@@ 6 Case Study
6.1 ReAgent Walk Through
To illustrate how recursive feedback enables robust multi-agent reasoning, we present a case where the system answers the question: “Which U.S. state has a capital city whose population is smaller than the state’s largest city, given that this state hosted the 1984 Summer Olympics?”, shown in Figure 6. The question is first decomposed by the Question Decomposer agent ($A_{Q}$) into four sub-questions: identifying the host state of the 1984 Olympics, retrieving its capital and largest city, comparing their populations, and returning the state if the capital is smaller. The Retriever agent ($A_{R}$) retrieves that California hosted the 1984 Olympics, with Sacramento as its capital and Los Angeles as its largest city. However, inconsistent population data for Sacramento (508k vs. 1.5M) triggers a local conflict, which is resolved by the Verifier agent ($A_{V}$) through local backtracking—discarding the unreliable 1.5M estimate. In this case, the local backtracking step is able to figure out the correct information, then the Answer Assembler ($A_{A}$) is able to gather all the information locally and then globally, and finally present the final answer to be California. We eliminate other parts when another conflict may occur, a full walk-through is presented in Appendix D. Another case study is presented in Appendix C
6.2 Comparison with Baseline Models
To elucidate differences in reasoning dynamics, we perform a case study comparing ReAgent with a single-agent reasoning baseline, O3 and an agentic baseline, Chain-of-Agents (CoA) Zhang et al. (2024) on the query ‘‘What is the name of the fight song of the university whose main campus is in Lawrence, Kansas and whose branch campuses are in the Kansas City metropolitan area?’’, as illustrated in Figure 7. The full comparison is presented in Appendix D. The example question requires two separate hops: (i) identifying which university satisfies the geographical constraints, and (ii) retrieving the “official” fight-song of that school. Although seemingly simple, the task hides a subtle distinction: the University of Kansas (KU) has both an official fight song (“Kansas Song”, adopted in 1928) and a far better-known rally tune (“I’m a Jayhawk”). Correctly answering therefore hinges on (a) verifying that KU is the only university matching the campus pattern, and (b) resolving the potential conflict between the two candidate songs.
The reasoning model O3 (left) performs the first hop successfully but then falls back on memorised popularity cues, assuming that the song most familiar to the public must be official.
The CoA (middle) multi-worker system executes three independent workers and lets a manager pick the majority vote. Worker 1 fails to locate any fight-song information due to the chunk division. Worker 2 repeats O3’s mistake. Worker 3 extracts the correct answer “Kansas Song” but discards it because it is less commonly used, and the third fails to locate any fight-song information. The manager simply adopts the majority answer (“I’m a Jayhawk”), which causes the error.
Our ReAgent (right) framework handles the same example with six specialized retriever agents answering the sub-questions generated by the Decomposer: (a) Q1–Q3 (entity verification). Three agents collectively confirm that only KU satisfies the main-campus/branch-campus constraints, eliminating alternative schools early. (b) Q4–Q5 (conflict surfacing). Independent retrieval agents surface both “I’m a Jayhawk” and “Kansas Song”, triggering an explicit conflict state. (c) Backtracking (Q6). The controller reverses to Q4 and re-weights evidence. The source for “Kansas Song” contains the enactment year (1928) and the keyword “official”, while “I’m a Jayhawk” is marked only as a “rally tune”. Therefore, our model successfully reach the correct answer “Kansas Song”. This demonstrates that, thanks to its conflict-detection and backtracking mechanisms, ReAgent can revisit earlier reasoning, resolve contradictory evidence, and consistently converge on the correct knowledge-grounded answer.
Figure 6: Case study: We illustrate the main steps of how ReAgent answers a question, where local backtracking is highlighted to resolve a conflict.